# On Erlang mixture approximations for differential equations

### v2

@@ 8.1 Modified logistic equation
Furthermore, we derive $\beta$ in (118) analytically.
We use the parameter values in Table 1, and the initial state is $x(t)=0.9$ for $t\leq t_{0}$. We use the same parameters as in Section 7.1 to identify Erlang mixture kernels of order $M=100$ in the bifurcation analysis with respect to the logistic growth rate, $\sigma$, and $M=350$ in the analysis with respect to the kernel parameter $\mu_{2}$. In Figure 5, the first row shows the eigenvalues of the Jacobian of the right-hand side function in the approximate system of ODEs (73). For completeness, we consider a very large range of values for $\sigma$, but for numerical reasons, we have not been able to investigate larger values of $\mu_{2}$. The second row shows the largest real part of the eigenvalues and when the parameters become sufficiently large, the steady state $\bar{x}=\kappa$ becomes unstable. The two bottom rows show simulations obtained with the numerical method for non-stiff DDEs for the parameter values indicated in the second row (the colors match). We use a memory horizon of $\Delta t_{h}=t_{f}-t_{0}=24$ mo and $10,000$ time steps. In both cases, the stability analysis based on the approximate set of ODEs (73) accurately predicts when the steady state is stable, marginally stable (where the largest real part of the eigenvalues is equal to zero), and unstable. For large values of $\sigma$, a limit cycle appears whereas for large values of $\mu_{2}$, the population density oscillates unstably. Figure 6 shows the identified kernels, the corresponding error, and the coefficients for different values of $\mu_{2}$. There is a similarity between the coefficients and the shape of the approximated kernel, e.g., in terms of bimodality. Furthermore, as $\mu_{2}$ increases, more sharp spikes occur in the coefficients.
@@
Figure 5: Bifurcation analysis with respect to the model parameter $\sigma$ (left column) and the kernel parameter $\mu_{2}$ (right column) for the modified logistic equation. First row: Eigenvalues. Second row: The largest real part of the eigenvalues. Third and bottom row: Simulations for selected parameter values (obtained with the numerical method described in Appendix B.1).
Figure 6: Top row: The true kernels and the corresponding Erlang mixture approximations for $M=100$ (left) and $M=350$ (right) obtained in connection with the bifurcation analysis shown in Figure 5. Middle row: The absolute errors of the kernel approximations. Bottom row: The coefficients in the Erlang mixture approximations. The colors are consistent across each column.
@@ 8.2 Nuclear fission
The kernels are described in Appendix C and represent 1) a nonuniform velocity profile in the external circulation of the molten salt (and neutron precursors) and 2) the (exponential) decay of the neutron precursors during the external circulation.
@@
We use $N=1000$ points and an optimality tolerance of $10^{-10}$ in fmincon to identify Erlang mixture kernel approximations of order $M=1000$. Furthermore, we use $\epsilon=10^{-13}$ and a tolerance of $10^{-14}$ in the bisection algorithm, we use Matlab’s integral with absolute and relative tolerances of $10^{-15}$ to approximate the integral in (118), and we use long double precision when evaluating the Erlang and Erlang mixture kernels. Figure 7 shows a Monte Carlo simulation for 1000 samples of $\kappa$ drawn from a normal distribution with mean $3\cdot 10^{-4}$ and standard deviation $7.5\cdot 10^{-5}$. For brevity of the presentation, we omit the plots of the neutron precursor group concentrations. The initial states are $C_{i}(t)=1$ kmol/cm3 for $i=1,\ldots,n$ and $\rho(t)=1.1\beta$ for $t\leq t_{0}$, and we use an absolute and relative tolerance of $10^{-8}$ in ode15s to simulate the approximate system of ODEs (73). Furthermore, we supply the analytical Jacobian of the right-hand side function, and we represent it as a sparse matrix. The figure also shows the relative difference between the simulations obtained 1) as described above and 2) with the numerical method for stiff DDEs described in Appendix B.2 for the mean value of $\kappa$. The latter method uses a time step size of $6.25\cdot 10^{-5}$ s, and the relative difference, $E_{r,i}:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$, in the $i$’th state variable and $n$’th time step is given by
@@
where $\hat{x}_{i}$ is the approximate solution to the ODEs (73) and $x_{i}$ is obtained with the numerical method for stiff DDEs. For each point in time, the 95% confidence intervals show the intervals from the 2.5 to the 97.5 percentiles and the span shows the interval of the minimum and maximum value of the states. For the neutron concentration, the confidence interval is almost symmetric (on the logarithmic axis) whereas the maximum is significantly higher above the mean than the minimum is below it. For the reactivity, the confidence interval and the span are hardly visible, and the simulation for the mean value of $\kappa$ is indistinguishable from the pointwise mean. The maximum relative difference for all eight state variables is $2.14\cdot 10^{-3}$. There are three main sources of error that affect this difference: 1) The kernel approximation error, 2) the error from using ode15s to simulate the approximate set of ODEs, and 3) the error of the numerical method described in Appendix B.2. If we use twice as large time steps in the latter, the maximum relative error is $4.28\cdot 10^{-3}$, i.e., twice as large (results not shown). Therefore, we believe that the maximum relative difference would be further reduced by using smaller time step sizes, i.e., that the main source of error is that of the numerical method for stiff DDEs and not the Erlang mixture kernel approximation. However, due to large memory requirements, it has not been possible to further reduce the time step size. Finally, Figure 8 shows the Erlang mixture approximations of the kernels and the corresponding errors. It also shows the sum of folded normal kernels in (156) for the values of $\mu_{1}$ and $\sigma_{1}$ shown in Table 2 and the coefficients in the Erlang mixture approximations. For some kernels, sharp spikes occur in the coefficients.
Figure 7: Monte Carlo simulation of the molten salt reactor model. Top row: The neutron concentration (left) and the reactivity (right). For each point in time, the span shows the interval of the minimum and maximum state, and the 95% confidence interval spans the 2.5 and 97.5 percentiles. The mean is computed pointwise, and the sample is the simulation corresponding to the mean value of $\kappa$. Bottom row: The pointwise relative error (137) between simulating the approximate ODEs (73) using ode15s and simulating the original DDEs using the numerical method for stiff DDEs described in Appendix B.2.
Figure 8: Left column: The Erlang mixture approximations of the kernels in (156) used in the Monte Carlo simulation shown in Figure 7 (top) and the corresponding approximation errors (bottom). Right column: The normalized sum of folded normal kernels in the expression (156) for the kernels (top), and the coefficients in the Erlang mixture kernel approximations. The colors are consistent across all but the top right figure.
@@ 9 Conclusions
In this paper, we propose an approximate approach for simulating and analyzing DDEs with distributed time delays based on conventional methods for ODEs. Specifically, we approximate the involved kernel by an Erlang mixture kernel and transform the resulting DDEs to ODEs using the LCT (Theorem 5.1). We prove that if the kernel $\alpha$ is regular (Definition 3.1) and the coefficients are chosen as in (17), then an infinite-order Erlang mixture approximation $\hat{\alpha}$ (Definition 4.2 where $M$ is infinite) converges pointwise to $\alpha$ (Theorem 4.26) and the integral of the absolute error converges to zero (Theorem 4.37) as the rate parameter, $a$, goes to infinity. Furthermore, we show that accurate stability analyses of the steady states of the original system of DDEs (2b)–(3) can be performed based on the approximate system of ODEs (73) if the kernel $\alpha$ is also exponentially bounded (Theorem 5.14). We also show that the solution to the approximate ODEs converges to that of the original system (Theorem 5.22). Additionally, we propose an approach for determining optimal Erlang mixture approximations based on bisection and least-squares estimation, and we use numerical examples to demonstrate that its accuracy and rate of convergence can be higher than using the theoretical expressions for the coefficients in (17). We also compare with the accuracy of an exponential mixture approximation from the literature. Finally, we present two numerical examples that demonstrate the efficacy of the proposed approach, and we compare the results with those obtained by two numerical methods (for non-stiff and stiff systems) formulated directly for DDEs with distributed time delays.
Future work may involve the derivation of error bounds, analysis of the error from truncating infinite-order Erlang mixture kernels (as in Section 6), an extension to Erlang mixture approximations with multiple rate parameters, and an extension to systems with time-varying kernels.

### v3

@@ 7.1 Modified logistic equation
Furthermore, we derive $\beta$ in Definition 3.1 analytically.
We use the parameter values in Table 2, and the initial state is $x(t)=0.9$ for $t\leq t_{0}$. We use the same parameters as in Section 6.1 to identify Erlang mixture kernels of order $M=100$ in the bifurcation analysis with respect to the logistic growth rate, $\sigma$, and $M=350$ in the analysis with respect to the kernel parameter $\mu_{2}$. Table 1 shows the resulting values of $t_{h}$ and $a$. In Figure 4, the first row shows the eigenvalues of the Jacobian of the right-hand side function in the approximate system of ODEs (19). For completeness, we consider a large range of values for $\sigma$, but for numerical reasons, we have not been able to investigate larger values of $\mu_{2}$. The second row shows the largest real part of the eigenvalues and when the parameters become sufficiently large, the steady state $\bar{x}=\kappa$ becomes unstable. The two bottom rows show simulations obtained with the numerical method for non-stiff DDEs for the parameter values indicated in the second row (the colors match). We use a memory horizon of $\Delta t_{h}=t_{f}-t_{0}=24$ mo and $10,000$ time steps. In both cases, the stability analysis based on the approximate set of ODEs (19) accurately predicts when the steady state is stable, marginally stable (where the largest real part of the eigenvalues is equal to zero), and unstable. For large values of $\sigma$, a limit cycle appears whereas for large values of $\mu_{2}$, the population density oscillates unstably. Figure 5 shows the identified kernels, the corresponding error, and the coefficients for different values of $\mu_{2}$. In the left column, the coefficients exhibit oscillation, whereas in the right column, there is a similarity between the coefficients and the shape of the approximated kernel, e.g., in terms of bimodality. The oscillations in the coefficients disappear for larger values of $M$, which also leads to larger values of $a$ (results not shown).
@@
Figure 4: Bifurcation analysis with respect to the model parameter $\sigma$ (left column) and the kernel parameter $\mu_{2}$ (right column) for the modified logistic equation. First row: Eigenvalues. Second row: The largest real part of the eigenvalues. Third and bottom row: Simulations for selected parameter values (obtained with the numerical method described in Appendix G.1).
Figure 5: Top row: The true kernels and the corresponding Erlang mixture approximations for $M=100$ (left) and $M=350$ (right) obtained in connection with the bifurcation analysis shown in Figure 4. Middle row: The absolute errors of the kernel approximations. Bottom row: The coefficients in the Erlang mixture approximations. The colors are consistent across each column, and the coefficients are visualized as solid lines to ease the interpretation of the results.
@@ 7.2 Nuclear fission
The kernels are described in Appendix H and represent 1) a nonuniform velocity profile in the external circulation of the molten salt (and neutron precursors) and 2) the (exponential) decay of the neutron precursors during the external circulation.
@@
We use $N=1000$ points to identify Erlang mixture kernel approximations of order $M=500$. Furthermore, we use $\epsilon=10^{-13}$ and a tolerance of $10^{-14}$ in the bisection algorithm (see Table 1), we use Matlab’s integral with absolute and relative tolerances of $10^{-15}$ to approximate $\beta$ in (5). Figure 6 shows a Monte Carlo simulation for 1000 samples of $\kappa$ drawn from a normal distribution with mean $3\cdot 10^{-4}$ and standard deviation $7.5\cdot 10^{-5}$. For brevity of the presentation, we omit the plots of the neutron precursor group concentrations. The initial states are $C_{i}(t)=1$ kmol/cm3 for $i=1,\ldots,n$ and $\rho(t)=1.1\beta$ for $t\leq t_{0}$, and we use an absolute and relative tolerance of $10^{-8}$ in ode15s to simulate the approximate system of ODEs (19). Furthermore, we supply the analytical Jacobian of the right-hand side function, and we represent it as a sparse matrix. The figure also shows the relative difference between the simulations obtained 1) as described above and 2) with the numerical method for stiff DDEs described in Appendix G.2 for the mean value of $\kappa$. The latter method uses a time step size of $6.25\cdot 10^{-5}$ s, and the relative difference, $E_{r,i}:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$, in the $i$’th state variable and $n$’th time step is given by
@@
where $\hat{x}_{i}$ is the approximate solution to the ODEs (19) and $x_{i}$ is obtained with the numerical method for stiff DDEs. For each point in time, the 95% confidence intervals show the intervals from the 2.5 to the 97.5 percentiles and the span shows the interval of the minimum and maximum value of the states. For the neutron concentration, the confidence interval is almost symmetric (on the logarithmic axis) whereas the maximum is significantly higher above the mean than the minimum is below it. For the reactivity, the confidence interval and the span are hardly visible, and the simulation for the mean value of $\kappa$ is indistinguishable from the pointwise mean. The maximum relative difference for all eight state variables is $2.14\cdot 10^{-3}$. There are three main sources of error that affect this difference: 1) The kernel approximation error, 2) the error from using ode15s to simulate the approximate set of ODEs, and 3) the error of the numerical method described in Appendix G.2. If we use twice as large time steps in the latter, the maximum relative error is $4.28\cdot 10^{-3}$, i.e., twice as large (results not shown). Therefore, we believe that the maximum relative difference would be further reduced by using smaller time step sizes, i.e., that the main source of error is that of the numerical method for stiff DDEs and not the Erlang mixture kernel approximation. However, due to large memory requirements, it has not been possible to further reduce the time step size. Finally, Figure 7 shows the Erlang mixture approximations of the kernels and the corresponding errors. It also shows the sum of folded normal kernels in (168) for the values of $\mu_{1}$ and $\sigma_{1}$ shown in Table 3 and the coefficients in the Erlang mixture approximations. The coefficients exhibit oscillations, which was also the case for the coefficients in the left column of Figure 5.
Figure 6: Monte Carlo simulation of the molten salt reactor model. Top row: The neutron concentration (left) and the reactivity (right). For each point in time, the span shows the interval of the minimum and maximum state, and the 95% confidence interval spans the 2.5 and 97.5 percentiles. The mean is computed pointwise, and the sample is the simulation corresponding to the mean value of $\kappa$. Bottom row: The pointwise relative error (90) between simulating the approximate ODEs (19) using ode15s and simulating the original DDEs using the numerical method for stiff DDEs described in Appendix G.2.
Figure 7: Left column: The Erlang mixture approximations of the kernels in (168) used in the Monte Carlo simulation shown in Figure 6 (top) and the corresponding approximation errors (bottom). Right column: The normalized sum of folded normal kernels in the expression (168) without the decay factor (top) and the coefficients in the Erlang mixture kernel approximations. The colors are consistent across all but the top right figure.
@@ 8 Conclusions
In this paper, we propose an approximate approach for simulating and analyzing DDEs with distributed time delays based on conventional methods for ODEs. Specifically, we approximate the involved kernel by an Erlang mixture kernel and transform the resulting DDEs to ODEs using the LCT. We prove that if the kernel $\alpha$ is regular (Definition 3.2) and the coefficients are chosen as in (15), then an Erlang mixture approximation, $\hat{\alpha}$, of order $M$ converges pointwise to $\alpha$ and the integral of the absolute error converges to zero as the rate parameter, $a$, and the order, $M$, go to infinity provided that the ratio $M/a$ goes to infinite as well (Proposition 4.2). The main theoretical result of the paper is Theorem 4.4, which states that 1) accurate stability analyses of the steady states of the original system of DDEs (3b)–(4) can be performed based on the approximate system of ODEs (19) if the kernel $\alpha$ is also exponentially bounded, and 2) the solution to the approximate ODEs converges uniformly to that of the original system. Additionally, we propose an approach for determining optimal Erlang mixture approximations based on bisection and least-squares estimation, and we use numerical examples to demonstrate that its accuracy and rate of convergence can be higher than using the theoretical expressions for the coefficients in (15). We also compare with the accuracy of an exponential mixture approximation from the literature. Finally, we present two numerical examples that demonstrate the efficacy of the proposed approach, and we compare the results with those obtained by two numerical methods (for non-stiff and stiff systems) formulated directly for DDEs with distributed time delays.
Future work may involve the derivation of error bounds, an extension to Erlang mixture approximations with multiple rate parameters, and an extension to systems with time-varying kernels.